# The formation of a soliton gas condensate for the focusing Nonlinear Schrödinger equation

### v1

@@ 6.4 Error function and global parametrix
$=\Lambda(z)\Big{(}e^{2\widetilde{y}(z)}\mathcal{Y}(z)\Big{)}^{\sigma_{3}/2}e^{-\ln(N)\frac{2c}{\tau}\int_{{\overline{A}}}^{z}\frac{{\rm d}s}{R(s)}\sigma_{3}}e^{i\pi\sigma_{3}/4}\,\Psi^{-1}\left(\varphi_{{\overline{A}}}(z)\right)\,e^{-i\pi\sigma_{3}/4}Q_{{\overline{A}}}^{-1}(z).$ (6.46)
Next we utilize the asymptotic behavior of the function $\Psi$ defined in (6.2), as presented in [28, p. 369]
@@
$J_{{\mathcal{E}}}(z)$ $=\frac{\sqrt{2}}{2}\,\Lambda(z)\Big{(}e^{2\widetilde{y}(z)}\mathcal{Y}(z)\Big{)}^{\sigma_{3}/2}e^{-\ln(N)\frac{2c}{\tau}\int_{{\overline{A}}}^{z}\frac{{\rm d}s}{R(s)}\sigma_{3}}e^{i\pi\sigma_{3}/4}e^{-2\left(\varphi_{{\overline{A}}}(z)\right)^{1/2}\sigma_{3}}$ (6.48)
$\left(I_{2}-\frac{1}{2}\begin{pmatrix}1&-i\\ -i&1\end{pmatrix}\begin{pmatrix}\mathcal{O}\left(\varphi_{{\overline{A}}}^{-1/2}(z)\right)&\mathcal{O}\left(\varphi_{{\overline{A}}}^{-1/2}(z)\right)\\ \\[0.5pt] \mathcal{O}\left(\varphi_{{\overline{A}}}^{-1/2}(z)\right)&\mathcal{O}\left(\varphi_{{\overline{A}}}^{-1/2}(z)\right)\end{pmatrix}+\ldots\right)\begin{pmatrix}1&-i\\ -i&1\end{pmatrix}\left(2\pi\varphi_{{\overline{A}}}^{1/2}(z)\right)^{\sigma_{3}/2}e^{-i\pi\sigma_{3}/4}Q_{{\overline{A}}}^{-1}(z)\,.$ (6.49)
Using (6.4), one can further simplify the previous expression because
$\exp\left(\left(-\ln(N)\frac{2c}{\tau}\int_{{\overline{A}}}^{z}\frac{{\rm d}s}{R(s)}+\frac{i\pi}{4}-2\varphi_{{\overline{A}}}^{1/2}(z)\right)\sigma_{3}\right)=e^{i\pi\sigma_{3}/4}.$ (6.50)
Therefore, $J_{\mathcal{E}}(z)$ becomes
@@
$\left(I_{2}-\frac{1}{2}\begin{pmatrix}1&-i\\ -i&1\end{pmatrix}\begin{pmatrix}\mathcal{O}\left(\varphi_{{\overline{A}}}^{-1/2}(z)\right)&\mathcal{O}\left(\varphi_{{\overline{A}}}^{-1/2}(z)\right)\\ \\[0.5pt] \mathcal{O}\left(\varphi_{{\overline{A}}}^{-1/2}(z)\right)&\mathcal{O}\left(\varphi_{{\overline{A}}}^{-1/2}(z)\right)\end{pmatrix}+\dots\right)\begin{pmatrix}1&-i\\ -i&1\end{pmatrix}\left(2\pi\varphi_{{\overline{A}}}^{1/2}(z)\right)^{\sigma_{3}/2}e^{-i\pi\sigma_{3}/4}Q_{{\overline{A}}}^{-1}(z)\,.$ (6.52)
@@
$Q_{{\overline{A}}}(z)=\frac{\sqrt{2}}{2}\Lambda(z)\Big{(}e^{2\widetilde{y}(z)}\mathcal{Y}(z)\Big{)}^{\sigma_{3}/2}e^{i\pi\sigma_{3}/4}\begin{pmatrix}1&-i\\ -i&1\end{pmatrix}\left(2\pi\left(\varphi_{{\overline{A}}}(z)\right)^{1/2}\right)^{\sigma_{3}/2}e^{-i\pi\sigma_{3}/4}.$ (6.53)
@@ 7 Solution to the NLS equation
$\tilde{A}(z)=\xi(z)\ e^{-\left(\log{N}g(\infty)+y(\infty)\right)\sigma_{3}}\,{\mathcal{E}}(z)\,e^{i\nu_{0}(z)\sigma_{3}}\Xi(z)e^{i\nu(z)\sigma_{3}}\,e^{\left(\log{N}g(z)+y(z)\right)\sigma_{3}}.$ (7.2)
@@
which, when combined with the definitions of $C(z)$, $\Lambda(z)$ and ${\mathcal{E}}(z)$ (see relations (3.3), (5.2) and (6.42)) leads to (7.2).
@@
Under the $N$-soliton condensate scattering data assumption, for all $(x,t)$ in a compact set $K$, there is $N_{0}$ so that for all $N>N_{0}$, $\psi_{SG}(x,t;N)$ satisfies the asymptotic description (7.8). Specifically, for $\left|\psi_{SG}\right|$, we have
$|\psi_{SG}(x,t;N)|^{2}=c^{2}\frac{\left(\Theta^{\prime}\left(\frac{\tau+1}{2};\tau\right)\right)^{2}}{\Theta(\tau;\tau)\Theta(0;\tau)}\frac{\Theta\left(-\frac{\zeta(x,N)}{2\pi}+\frac{1-\tau}{2};\tau\right)\Theta\left(-\frac{\zeta(x,N)}{2\pi}+\frac{\tau-1}{2};\tau\right)}{\Theta^{2}\left(-\frac{\zeta(x,N)}{2\pi};\tau\right)}+\mathcal{O}\left(\frac{1}{\log{N}}\right),$ (7.4)
where the error term $\mathcal{O}\left(\frac{1}{\log{N}}\right)$ is uniform for all $(x,t)$ in the compact set $K$. Here $\Theta(z;\tau)$ is the Jacobi Theta-3 function as defined in [12]. The constants $\tau,c$ and $\zeta(x,N)$ are given by
$c=\frac{1}{2}\left(\int_{A}^{-{\overline{A}}}\frac{1}{R_{+}(s)}{\rm d}s\right)^{-1}\,,\quad\tau=2c\int_{-{\overline{A}}}^{-A}\frac{1}{R(s)}{\rm d}s\,,$ (7.5)
$\zeta(x,N)=-\frac{\tau\Delta(x)-2\ln(N)}{2c\pi}\left(\int_{-{\overline{A}}}^{-A}\frac{s^{2}}{R(s)}{\rm d}s-\tau\int_{A}^{-{\overline{A}}}\frac{s^{2}}{R_{+}(s)}{\rm d}s\right),$ (7.6)
where $R(z)=(z-A)^{\frac{1}{2}}(z+A)^{\frac{1}{2}}(z-{\overline{A}})^{\frac{1}{2}}(z+{\overline{A}})^{\frac{1}{2}}$, the branch-cuts are the canonical ones, and the integration is performed on the straight line connecting the endpoints. The function $\Delta(x)$ is defined in terms of the functions $\rho(z)$, $h(z)$ as follows
$\Delta(x)=\left(\int_{-A}^{-{\overline{A}}}\frac{1}{R(s)}{\rm d}s\right)^{-1}\left(-\int_{A}^{-{\overline{A}}}\frac{\log\,(2\pi h(s)\rho(s))+2isx}{R_{+}(s)}{\rm d}s+\int_{-A}^{{\overline{A}}}\frac{\log\left(2\pi h(s)\rho(s)\right)-2isx}{R_{+}(s)}{\rm d}s\right)\,.$ (7.7)
@@
$\begin{split}\psi_{SG}(x,t;N)&=\lim_{z\to\infty}2iz\,\xi(z)\,s_{12}(z;x,t)\,\exp\left(-\ln(N)(g(\infty)+g(z))+i\nu_{0}-i\nu(z)-y(z)-y(\infty)\right)+\mathcal{O}\left(\frac{1}{\log{N}}\right)\\ {\overline{\psi_{SG}(x,t;N)}}&=\lim_{z\to\infty}2iz\,\xi(z)\,s_{21}(z;x,t)\,\exp\left(+\ln(N)(g(\infty)+g(z))-i\nu_{0}+i\nu(z)+y(z)+y(\infty)\right)+\mathcal{O}\left(\frac{1}{\log{N}}\right)\end{split}$ (7.8)
Therefore, since the two previous limits are finite, we can compute the modulus of the solution as
$\begin{split}|\psi_{SG}(x,t;N)|^{2}=\lim_{z\to\infty}&-\frac{4z^{2}\xi^{2}(z)}{F_{1}(\infty)F_{2}(\infty)}\frac{\Theta\left(-\mathcal{A}(z)+\frac{1+3\tau}{4};\tau\right)\Theta\left(-\mathcal{A}(z)-\frac{\zeta(x,N)}{2\pi}+\frac{1-\tau}{4};\tau\right)}{\Theta\left(-\mathcal{A}(z)+\frac{1}{2};\tau\right)\Theta\left(-\mathcal{A}(z)+\frac{\tau}{2};\tau\right)}\\ &\times\frac{\Theta\left(\mathcal{A}(z)+\frac{3+\tau}{4};\tau\right)\Theta\left(\mathcal{A}(z)-\frac{\zeta(x,N)}{2\pi}+\frac{\tau-1}{4};\tau\right)}{\Theta\left(\mathcal{A}(z)+\frac{1}{2};\tau\right)\Theta\left(\mathcal{A}(z)+\frac{\tau}{2};\tau\right)}+\mathcal{O}\left(\frac{1}{\log{N}}\right)\,,\end{split}$ (7.9)
where we used the definition of $s_{12},s_{21}$ in (5.50). We notice that the terms in the previous equations have the following asymptotics for large-$z$
$\xi(z)\sim 1+O(z^{-1})$ (7.10)
$\frac{\Theta\left(-\mathcal{A}(z)+\frac{1+3\tau}{4};\tau\right)\Theta\left(-\mathcal{A}(z)-\frac{\zeta(x,N)}{2\pi}+\frac{1-\tau}{4};\tau\right)}{\Theta\left(-\mathcal{A}(z)+\frac{1}{2};\tau\right)\Theta\left(-\mathcal{A}(z)+\frac{\tau}{2};\tau\right)}\sim\frac{\theta_{1}(N)}{z}+O(z^{-2})$ (7.11)
$\frac{\Theta\left(\mathcal{A}(z)+\frac{3+\tau}{4};\tau\right)\Theta\left(\mathcal{A}(z)-\frac{\zeta(x,N)}{2\pi}+\frac{\tau-1}{4};\tau\right)}{\Theta\left(\mathcal{A}(z)+\frac{1}{2};\tau\right)\Theta\left(\mathcal{A}(z)+\frac{\tau}{2};\tau\right)}\sim\frac{\theta_{2}(N)}{z}+O(z^{-2})$ (7.12)
where $\theta_{1}(N),\theta_{2}(N)$ are defined as
$\begin{split}\theta_{1}(N)&=-c\,\Theta^{\prime}\left(\frac{\tau+1}{2};\tau\right)\frac{\Theta\left(-\frac{\zeta(x,N)}{2\pi}+\frac{1-\tau}{2};\tau\right)}{F_{1}(\infty)\Theta\left(\frac{3-\tau}{4};\tau\right)\Theta\left(\frac{1+\tau}{4};\tau\right)}\\ \theta_{2}(N)&=c\,\Theta^{\prime}\left(\frac{\tau+1}{2};\tau\right)\frac{\Theta\left(-\frac{\zeta(x,N)}{2\pi}+\frac{\tau-1}{2};\tau\right)}{F_{2}(\infty)\Theta\left(\frac{1+\tau}{4};\tau\right)\Theta\left(\frac{3\tau-1}{4};\tau\right)}\,,\end{split}$ (7.13)
and where the constant $c\in{\mathbb{R}}_{-}$ was previously computed is independent of $N$. In this notation, we can rewrite (7.9) as
$|\psi_{SG}(x,t;N)|^{2}=c^{2}\frac{\left(\Theta^{\prime}\left(\frac{\tau+1}{2};\tau\right)\right)^{2}}{\Theta(\tau;\tau)\Theta(0;\tau)}\frac{\Theta\left(-\frac{\zeta(x,N)}{2\pi}+\frac{1-\tau}{2};\tau\right)\Theta\left(-\frac{\zeta(x,N)}{2\pi}+\frac{\tau-1}{2};\tau\right)}{\Theta^{2}\left(-\frac{\zeta(x,N)}{2\pi};\tau\right)}+\mathcal{O}\left(\frac{1}{\log{N}}\right)$ (7.14)
where we also used eqs. (5.48)-(5.49). ∎

### v2

@@ 6.4 Error function and global parametrix
$=\Lambda(z)\Big{(}e^{2\widetilde{y}(z)}\mathcal{Y}(z)\Big{)}^{\sigma_{3}/2}e^{2\varphi_{{\overline{A}}}(z)^{1/2}\sigma_{3}}e^{i\pi\sigma_{3}/4}\,\Psi^{-1}\left(\varphi_{{\overline{A}}}(z)\right)\,e^{-i\pi\sigma_{3}/4}Q_{{\overline{A}}}^{-1}(z).$ (6.38)
Next we utilize the asymptotic behavior of the function $\Psi$ defined in (6.2), as presented in [29, p. 369]
@@
$J_{{\mathcal{E}}}(z)$ $=\frac{\sqrt{2}}{2}\,\Lambda(z)\Big{(}e^{2\widetilde{y}(z)}\mathcal{Y}(z)\Big{)}^{\sigma_{3}/2}e^{2\varphi_{{\overline{A}}}(z)^{1/2}\sigma_{3}}e^{i\pi\sigma_{3}/4}e^{-2\varphi_{{\overline{A}}}(z)^{1/2}\sigma_{3}}$ (6.40)
$\left(I-\frac{1}{2}\begin{pmatrix}1&-i\\ -i&1\end{pmatrix}\begin{pmatrix}\mathcal{O}\left(\varphi_{{\overline{A}}}^{-1/2}(z)\right)&\mathcal{O}\left(\varphi_{{\overline{A}}}^{-1/2}(z)\right)\\ \\[0.5pt] \mathcal{O}\left(\varphi_{{\overline{A}}}^{-1/2}(z)\right)&\mathcal{O}\left(\varphi_{{\overline{A}}}^{-1/2}(z)\right)\end{pmatrix}+\ldots\right)\begin{pmatrix}1&-i\\ -i&1\end{pmatrix}\left(2\pi\varphi_{{\overline{A}}}^{1/2}(z)\right)^{\sigma_{3}/2}e^{-i\pi\sigma_{3}/4}Q_{{\overline{A}}}^{-1}(z)\,.$
The last one simplifies to:
@@
$\left(I-\frac{1}{2}\begin{pmatrix}1&-i\\ -i&1\end{pmatrix}\begin{pmatrix}\mathcal{O}\left(\varphi_{{\overline{A}}}^{-1/2}(z)\right)&\mathcal{O}\left(\varphi_{{\overline{A}}}^{-1/2}(z)\right)\\ \\[0.5pt] \mathcal{O}\left(\varphi_{{\overline{A}}}^{-1/2}(z)\right)&\mathcal{O}\left(\varphi_{{\overline{A}}}^{-1/2}(z)\right)\end{pmatrix}+\dots\right)\begin{pmatrix}1&-i\\ -i&1\end{pmatrix}\left(2\pi\varphi_{{\overline{A}}}^{1/2}(z)\right)^{\sigma_{3}/2}e^{-i\pi\sigma_{3}/4}Q_{{\overline{A}}}^{-1}(z)\,.$ (6.42)
@@
$Q_{{\overline{A}}}(z)=\frac{\sqrt{2}}{2}\Lambda(z)\Big{(}e^{2\widetilde{y}(z)}\mathcal{Y}(z)\Big{)}^{\sigma_{3}/2}e^{i\pi\sigma_{3}/4}\begin{pmatrix}1&-i\\ -i&1\end{pmatrix}\left(2\pi\varphi_{{\overline{A}}}(z)^{1/2}\right)^{\sigma_{3}/2}e^{-i\pi\sigma_{3}/4}.$ (6.43)
@@ 7 Solution to the NLS equation
$\tilde{A}(z)=\xi(z)\ e^{-\left(\log{N}g_{\infty}+y_{\infty}\right)\sigma_{3}}\,\,e^{i\nu_{0}(z)\sigma_{3}}\Xi(z)e^{i\nu(z)\sigma_{3}}\,e^{\left(\log{N}g(z)+y(z)\right)\sigma_{3}}.$ (7.2)
@@
which, when combined with the definitions of $C(z)$, $\Lambda(z)$ (see relations (3.3), (5.2) and (6.34)) leads to (7.2).
@@
Under the $N$-soliton condensate scattering data assumption, for all $(x,t)$ in a compact set ${\mathfrak{K}}$, there is $N_{0}$ so that for all $N>N_{0}$, $\psi_{SG}(x,t;N)$ has the following asymptotic behavior
$\begin{split}\psi_{SG}&(x,t;N)\\ &=-2i\frac{\Re(A)\Im(A)}{|A|}\operatorname{sd}\left(2|A|x+\frac{2|A|}{\pi}\Re\left(\int_{\eta_{1}}\frac{\log(2\pi h(s)\rho(s))}{R_{+}(s)}{\rm d}s\right);\sin(\theta)\right)e^{it(A^{2}+{\overline{A}}^{2})+i\phi_{0}}+\mathcal{O}\left(\frac{1}{\log{N}}\right),\end{split}$ (7.4)
where the error term $\mathcal{O}\left(\frac{1}{\log{N}}\right)$ is uniform for all $(x,t)$ in the compact set ${\mathfrak{K}}$. Here $\operatorname{sd}$ is the Jacobi elliptic function $\operatorname{sd}$ [13, Chap. 22], and $R(z),\phi_{0}$ are given by
$R(z)=(z-A)^{\frac{1}{2}}(z+A)^{\frac{1}{2}}(z-{\overline{A}})^{\frac{1}{2}}(z+{\overline{A}})^{\frac{1}{2}}\,,$ (7.5)
$\phi_{0}=\ln(N)\frac{K(\cos(\theta))}{K(\sin(\theta))}+\Re\left(\int_{\eta_{1}}\frac{s\ln(2\pi h(s)\rho(s))}{R_{+}(s)}{\rm d}s\right)\,,\quad\theta=\arg(A)$ (7.6)
where the function $K(k)$ is the complete elliptic integral of the fist kind [13, Chap. 19], the branch-cuts are the canonical ones, and the integration is performed on the straight line connecting the endpoints.
@@
$\begin{split}\psi_{SG}(x,t;N)&=\lim_{z\to\infty}2iz\,\xi(z)\,s_{12}(z;x,t)\,\exp\left(-\ln(N)(g_{\infty}+g(z))+i\nu_{0}-i\nu(z)-y(z)-y_{\infty}\right)+\mathcal{O}\left(\frac{1}{\log{N}}\right)\\ &=\lim_{z\to\infty}2iz\,\xi(z)\,\Pi(z)\,\exp\left(-\ln(N)(g_{\infty}+g(z))+i\nu_{0}-i\nu(z)-y(z)-y_{\infty}+i\delta(z)\right)+\mathcal{O}\left(\frac{1}{\log{N}}\right)\,,\\ &=2i\left(\lim_{z\to\infty}z\Pi(z)\right)\exp(-2\ln(N)g_{\infty}-2y_{\infty}+2i\nu_{0})+\mathcal{O}\left(\frac{1}{\log{N}}\right)\end{split}$ (7.7)
where we set
$\Pi(z)=\frac{1}{Z_{2}}\frac{\Theta\left(\mathcal{A}(z)+\frac{3+\tau}{4};\tau\right)\Theta\left(\mathcal{A}(z)-\frac{\zeta(x,N)}{2\pi}+\frac{\tau-1}{4};\tau\right)}{\Theta\left(\mathcal{A}(z)+\frac{1}{2};\tau\right)\Theta\left(\mathcal{A}(z)+\frac{\tau}{2};\tau\right)}\,.$ (7.8)
Focusing on the exponential term, one can compute the terms $g_{\infty},y_{\infty},\nu_{0}$ from Proposition 3.6, Proposition 3.7 and Proposition 5.5 obtaining that
$g_{\infty}$ $=\frac{1}{2\tau}$ (7.9)
$y_{\infty}$ $=\frac{it}{2}(A^{2}+{\overline{A}}^{2})+\frac{c}{\tau}\left(\pi x+\Re\left(\int_{\eta_{1}}\frac{\log\,(2\pi h(s)\rho(s))}{R_{+}(s)}{\rm d}s\right)\right)+\frac{i}{\pi}\Re\left(\int_{\eta_{1}}\frac{s\log(2\pi h(s)\rho(s))}{R_{+}(s)}{\rm d}s\right)$ (7.10)
$\nu_{0}$ $=i\left(4c\pi x+2\ln(N)+4c\Re\left(\int_{\eta_{1}}\frac{\log(2\pi h(s)\rho(s))}{R_{+}(s)}{\rm d}s\right)\right)\left(\frac{1}{4}+\frac{1}{4\tau}\right)\,.$ (7.11)
With tears, one deduces that
$\exp(-2\ln(N)g_{\infty}-2y_{\infty}+2i\nu_{0})=\exp\left(\frac{i\zeta(x,N)}{2}+it(A^{2}+{\overline{A}}^{2})+i\phi_{0}\right)\,,$ (7.12)
where $\phi_{0}$ is defined in (7.5) and $\zeta(x,N)$ in (5.21). We now consider
$2i\lim_{z\to\infty}z\Pi(z)=-2ic\frac{\Theta^{\prime}\left(\frac{\tau+1}{2}\right)\Theta\left(-\frac{\zeta(x,N)}{2\pi}+\frac{\tau-1}{2}\right)\Theta\left(\frac{3-\tau}{4}\right)}{\Theta(0)\Theta\left(\frac{3\tau-1}{4}\right)\Theta\left(\frac{\zeta(x,N)}{2\pi}\right)}\,,$
where we used Proposition 5.6. We can further simplify the previous equation by expressing it first in terms of the Jacobi theta functions [13, Chap. 20], and then in terms of the $\operatorname{sd}$ elliptic function [13, Chap. 22] as
$2i\lim_{z\to\infty}z\Pi(z)=-2i\frac{\Re(A)\Im(A)}{|A|}\operatorname{sd}\left(2|A|x+\frac{2|A|}{\pi}\Re\left(\int_{\eta_{1}}\frac{\log(2\pi h(s)\rho(s))}{R_{+}(s)}{\rm d}s\right);\sin(\theta)\right)e^{-i\frac{\zeta(x,N)}{2}}\,.$ (7.13)
Combing (7.12)-(7.13), we deduce the result.
@@
8 Connection to the kinetic theory
As discussed in Section 2.2, in order to connect to the kinetic theory of solitons, we will consider a tracer soliton interacting with a sequence of $N$-soliton condensate solutions. That is, we consider RHP 2.1, with an additional pair of poles $\{\lambda_{o},{\overline{\lambda}}_{o}\}$, where $\lambda_{o}$ is a point in the discrete spectrum in $\mathbb{C}_{+}$ and outside the curve $\Gamma_{1}$. By symmetry ${\overline{\lambda}}_{o}$ is its conjugate in $\mathbb{C}_{-}$ and outside the curve $\Gamma_{2}$. Specifically, RHP 2.1, for a pure solitonic solution, becomes:
RHP 8.1.
Given spectral data $\mathcal{D}(t)$, find a $2\times 2$ matrix-valued function $M(z)$ such that:
  • $\ {M}(z)$ is meromorphic in ${\mathbb{C}}$, with simple poles at the points $\{\lambda_{j},\overline{\lambda_{j}}\}_{j=1}^{N}$ and the pair $\{\lambda_{o},{\overline{\lambda}}_{o}\}$.
  • At each of the poles $\lambda_{j}$ in $\mathbb{C}_{+}$, $\ {M}(z)$ satisfies the residue condition
    $\operatorname*{\textrm{res}}_{\lambda_{j}}M(z)=\lim_{z\to\lambda_{j}}\left[M(z)\begin{pmatrix}0&0\\ c_{j}e^{2i\theta(\lambda_{j};x,t)}&0\end{pmatrix}\right]\,,$ (8.1)
    and at each pole $\overline{\lambda_{j}}$ in $\mathbb{C}_{-}$:
    $\operatorname*{\textrm{res}}_{{\overline{\lambda_{j}}}}M(z)=\lim_{z\to{\overline{\lambda}}_{j}}\left[M(z)\begin{pmatrix}0&-{\overline{c}}_{j}e^{-2i\theta({\overline{\lambda}}_{j};x,t)}\\ 0&0\end{pmatrix}\right].$ (8.2)
  • At $z=\lambda_{o}$, $M(z)$ satisfies the residue condition:
    $\operatorname*{\textrm{res}}_{\lambda_{o}}M(z)=\lim_{z\to\lambda_{o}}\left[M(z)\begin{pmatrix}0&0\\ c_{o}e^{2i\theta(\lambda_{o};x,t)}&0\end{pmatrix}\right]\,,$ (8.3)
    and at $z=-{\overline{\lambda}}_{o}$, we have:
    $\operatorname*{\textrm{res}}_{{\overline{\lambda_{o}}}}M(z)=\lim_{z\to{\overline{\lambda}}_{o}}\left[M(z)\begin{pmatrix}0&-{\overline{c}}_{o}e^{-2i\theta({\overline{\lambda}}_{o};x,t)}\\ 0&0\end{pmatrix}\right].$ (8.4)
  • $\ {M}(z)=\ I+O(z^{-1})$, as $z\to\infty$.
The analysis described in the previous sections still applies, except that we need to account at each step for the poles $z=\lambda_{o},{\overline{\lambda}}_{o}$. So the transformations of Section 3, encircling the poles $\lambda_{j}$ with $\Gamma_{1}$ (and similarly ${\overline{\lambda}}_{j}$ with $\Gamma_{2}$), and replacing the residue conditions at each pole with a jump across the contours $\Gamma_{1}$ and $\Gamma_{2}$, which we then collapse onto $\eta_{1}$ and $\eta_{2}$ are carried out in exactly the same way, but ensuring that $\lambda_{o}$ and ${\overline{\lambda}}_{o}$ are outside the contours. This implies that the series of transformations:
$M(z)\to A(z)\to\widetilde{A}(z)\to B(z)$
all apply. Since $\lambda_{o}$ and ${\overline{\lambda_{o}}}$ are outside $\Gamma_{1}\cup\Gamma_{2}$, the residue conditions at $\lambda_{o}$ (8.3) and ${\overline{\lambda_{o}}}$ (8.4) may be rewritten, replacing $M$ with $B$:
$\operatorname*{\textrm{res}}_{\lambda_{o}}B(z)=\lim_{z\to\lambda_{o}}\left[B(z)\begin{pmatrix}0&0\\ c_{o}e^{2i\theta(\lambda_{o};x,t)}&0\end{pmatrix}\right]\,,$ (8.5)
and at $z=-{\overline{\lambda}}_{o}$, we have:
$\operatorname*{\textrm{res}}_{{\overline{\lambda_{o}}}}B(z)=\lim_{z\to{\overline{\lambda}}_{o}}\left[B(z)\begin{pmatrix}0&-{\overline{c}}_{o}e^{-2i\theta({\overline{\lambda}}_{o};x,t)}\\ 0&0\end{pmatrix}\right].$ (8.6)
The next transformation introduces the function $C(z)$, which preserves its jump but alters the residue condition at $z=\lambda_{o},{\overline{\lambda}}_{o}$. Specifically, $C(z)$ satisfies the following RHP:
RHP 8.2.
We look for a function $C(z)\in\text{Mat}({\mathbb{C}},2)$ such that:
  • The boundary values satisfy $C_{+}(z)=C_{-}(z)J_{C}(z)$, where the jump matrix $J_{C}(z)$ is given as in RHP 3.2.
  • At $z=\lambda_{o}$, $C(z)$ satisfies the residue condition:
    $\operatorname*{\textrm{res}}_{\lambda_{o}}C(z)=\lim_{z\to\lambda_{o}}\left[C(z)\begin{pmatrix}0&0\\ c_{o}e^{2i\theta(\lambda_{o};x,t)}e^{-2\left(\ln(N)g(\lambda_{o})+y(\lambda_{o})\right)}&0\end{pmatrix}\right]$ (8.7)
    and at $z={\overline{\lambda}}_{o}$, $C(z)$ satisfies the residue condition:
    $\operatorname*{\textrm{res}}_{{\overline{\lambda}}_{o}}C(z)=\lim_{z\to{\overline{\lambda}}_{o}}\left[C(z)\begin{pmatrix}0&-{\overline{c_{o}}}e^{-2i\theta({\overline{\lambda}}_{o};x,t)}e^{2\left(\ln(N)g({\overline{\lambda}}_{o})+y({\overline{\lambda}}_{o})\right)}\\ 0&0\end{pmatrix}\right].$ (8.8)
  • $C(z)=I+\mathcal{O}(1/z)$, as $z\to\infty$.
The analysis carried out in Section 4 may then be repeated, from $C$ to $D$ and the lens-opening from $D$ to $E$, so long as the additional poles $\lambda_{o}$ and ${\overline{\lambda_{o}}}$ are outside of the domains where transformations are made. Next, the global approximation constructed in Section 5 is replaced by a modified global approximation that includes the additional poles $\lambda_{o}$ and $\bar{\lambda_{o}}$. The quantity $\Lambda$ solving Riemann-Hilbert problem 5.1 then also has these two simple poles, and satisfies the residue conditions
$\operatorname*{\textrm{res}}_{\lambda_{o}}\Lambda(z)=\lim_{z\to\lambda_{o}}\left[\Lambda(z)\begin{pmatrix}0&0\\ c_{o}e^{2i\theta(\lambda_{o};x,t)}e^{-2\left(\ln(N)g(\lambda_{o})+y(\lambda_{o})\right)}&0\end{pmatrix}\right]$ (8.9)
and
$\operatorname*{\textrm{res}}_{{\overline{\lambda}}_{o}}\Lambda(z)=\lim_{z\to{\overline{\lambda}}_{o}}\left[\Lambda(z)\begin{pmatrix}0&-{\overline{c_{o}}}e^{-2i\theta({\overline{\lambda}}_{o};x,t)}e^{2\left(\ln(N)g({\overline{\lambda}}_{o})+y({\overline{\lambda}}_{o})\right)}\\ 0&0\end{pmatrix}\right].$ (8.10)
and $\Lambda(z)=I+\mathcal{O}(1/z)$, as $z\to\infty$.
The local approximations as constructed in Section 6 may be used without any changes, and so we arrive at an analogous Riemann-Hilbert problem for the error matrix $\mathcal{E}$, which is again in the small norm setting. At the end of all of this, the previously determined $\psi_{SG}(x,t;N)$ (the soliton gas condensate solution) whose asymptotics are determined from the Riemann-Hilbert problem for $\Lambda$, is more complicated, and we use $\psi_{SG+T}(x,t;\lambda_{o},N)$ to denote this new solution. It describes the leading order asymptotic interaction between the tracer soliton determined by $\lambda_{o}$ and $c_{o}$ and the condensate gas. It is straightforward to follow the outline presented above in this section and establish that
$\psi_{N+1}(x,t;\lambda_{o})=\psi_{SG+T}(x,t;\lambda_{o},N)+\mathcal{O}\left(\frac{1}{N}\right),$ (8.11)
with error that is uniformly controlled on any compact domain in $(x,t)$.
Heuristically, the tracer soliton is localized within the region where the residue condition is $\mathcal{O}(1)$ (i.e., neither large nor small). To achieve this, we choose $c_{o}$ proportional to $\exp{(2\left(\ln(N)\Re[g(\lambda_{o})]\right)}$ so that the residue condition for the tracer soliton is $\mathcal{O}(1)$ at $t=0$ and $x$ near 00. As $t$ evolves, the tracer soliton’s “effective position” is approximated by the equation
$\Re[2i\theta({\lambda}_{o};x,t)]=0$
which then leads to the equation
$x(t)=-4\Re[\lambda_{o}]t.$ (8.12)
The analysis outlined above shows that the velocity of the tracer soliton is effectively constant, and the same as if the soliton were propagating in the absence of the background condensate.
Remark 8.1.
It would be possible to carry out the more precise analysis in Section 6 of [20] and prove that the global maximum of $\psi_{N+1}(x,t;\lambda_{o})$ evolves in an oscillatory fashion, with average velocity $-4\Re[\lambda_{0}]$. Since the details of this analysis are similar to those in [20], we omit them.
In 2003 [16], El derived a system of kinetic equations for the velocity $v(x,t,z)$ of a trial soliton propagating through a dense KdV soliton gas:
$v(z)=4z^{2}+\frac{1}{z}\int_{0}^{\infty}\ln\left|\frac{s+z}{s-z}\right|(v(z)-v(s))f(s){\rm d}s$ (8.13)
where the local soliton density $f(x,t,z)$ satisfies the continuity equation
$f_{t}+(vf)_{x}=0.$
For the $N$-soliton consdensate solutions of the NLS equation considered here, our analysis leads to an analogous integral equation:
$v(z)=-4\Re[z]+\frac{1}{z}\int_{0}^{\infty}\ln\left|\frac{s+z}{s-z}\right|(v(z)-v(s))f(s){\rm d}s.$ (8.14)
In this setting, $f$ is linked to the $x$-derivative of the jump of $g^{\prime}$ across $\eta_{3}$. While the jump of $g^{\prime}$ across $\eta_{3}$ is non-trivial, it is actually independent of $x$, and so $f=0$. This not only satisfies the continuity equation trivially but also simplifies equation (8.14) to $v(z)=-4\Re[z]$.
In other words, the velocity equation (8.12) for the tracer soliton obtained through our analysis provides the first rigorous justification of the kinetic theory, in which a deterministic sequence of $N+1$ soliton solutions depicts the dual interpretation of a tracer soliton interacting with a gas comprised of a finite but asymptotically large collection of solitons as they form a soliton condensate gas.